# HIGHT: Hierarchical Graph Tokenization for Graph-Language Alignment

### v1

@@ 3.1 Node-Centric Tokenization
$p_{\mathbf{\theta}}({\bm{a}}|{\bm{q}},{\bm{h}})=\Pi_{i=1}^{l_{a}}p_{\mathbf{\theta}}({\bm{a}}_{i}|{\bm{q}},f_{n}({\bm{h}}_{1}),...,f_{n}({\bm{h}}_{n}),{\bm{a}}_{<i}),$ (4)
where ${\bm{h}}_{1},...,{\bm{h}}_{n}$ are node embeddings from a GNN typically pretrained through self-supervised learning on large-scale molecular datasets such as ZINC250k (Sterling and Irwin, 2015), $f_{n}$ is the corresponding adapter to align the node tokens to the LLM tokens. There are various options for tokenizing a molecule, given that a simple supervised trained GNN could produce meaningful tokens (Liu et al., 2023e). In this work, we consider a state-of-the-art node-centric tokenizer from Mole-BERT (Xia et al., 2023) that pretrains a VQVAE (van den Oord et al., 2017) with masked atoms modeling. It constructs a codebook ${\mathcal{Z}}$ to discretize atoms:
$z_{u}=\operatorname*{arg\,min}_{i}||{\bm{h}}_{u}-{\bm{e}}_{i}||_{2},$ (5)
where $z_{u}\in{\mathcal{Z}}$ is the quantized index of atom $u$, and ${\bm{e}}_{v}$ is the codebook embedding of the $i$-th entry. The codebook is trained through a reconstruction loss with respect to some attribute ${\bm{v}}_{i}$ of atom $i$:
${\mathcal{L}}_{r}=\frac{1}{n}\sum_{i=1}^{n}(1-\frac{{\bm{v}}_{i}^{T}\hat{{\bm{v}}_{i}}}{||{\bm{v}}_{i}||\cdot||\hat{{\bm{v}}_{i}}||})^{\gamma}+\frac{1}{n}\sum_{i=1}^{n}||\text{sg}[{\bm{h}}_{i}]-{\bm{e}}_{z_{i}}||_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{n}||\text{sg}[{\bm{e}}_{z_{i}}]-{\bm{h}}_{i}||_{2}^{2},$ (6)
where $\text{sg}[\cdot]$ is the stop-gradient operator in the straight-through estimator (Bengio et al., 2013), $\hat{{\bm{v}}_{i}}$ is the reconstructed attribute of atom $i$ with a decoder, and $\beta$ is a hyperparamter. In Mole-BERT, the attribute is simply the type of atom masked during training. Mole-BERT also manually partitions the codebook into groups of common atoms such as carbon, nitrogen, and oxygen in order to avoid codebook conflicts (Xia et al., 2023).
@@
Figure 2: Illustration of hallucination caused by node-centric tokenization. With only node-level tokens (i.e., discrete atom embeddings), LLMs have to identify and connect the nodes within a specific functional group in order to align useful molecular structures in a molecule to the corresponding language descriptions. Yet, due to the arbitrary order of atoms and position biases in LLMs, it is harder to distinguish each functional group, which further leads to hallucination and subpar alignment.
Intuitively, the trained atom tokens encode some contextual information, such as the neighbors of the atoms. However, node-centric tokenization makes the molecule-language alignment more challenging, as LLMs have to additionally find the specific nodes to align the corresponding texts during the instruction tuning process. It often encounters the underspecification issue during the alignment process. For example, in molecules, motifs or functional groups usually capture rich semantics, and often share many common atoms such as carbon, nitrogen, and oxygen (Bohacek et al., 1996). As shown in Fig. 2, given the input molecule with SMILES “C1=CC(C(C(=C1)C(=O)O)O)O”, LGLMs with node-centric tokenization can struggle in distinguishing the carboxylic acid (“R-COOH”) from the hydroperoxide (“R-OOH”). In fact, both functional groups contain two oxygen atoms and a hydrogen atom while the benzene ring will also provide additional carbon atoms to the hydroperoxide group. Merely looking into the atoms makes it hard for LGLMs to distinguish which functional group is present in the molecule. Furthermore, due to the loss of positional information in the node-centric tokenization (Liang et al., 2023; Cao et al., 2023), the limited expressivity of GNNs (Xu et al., 2019) and the positional biases of auto-regressive LLMs (Lu et al., 2022), it is more challenging for the inner LLM to relate the node-level information within a motif, which will lead to more serious performance degeneration of the graph-language alignment.
@@ 3.2 Motif Hallucination
To understand the issue of node-centric tokenization more clearly, we construct a simple benchmark called MotifHallu, which measures the hallucination of common functional groups by LGLMs. Specifically, we consider the $38$ common functional groups in RDKit [^1] https://github.com/rdkit/rdkit/blob/master/Data/FunctionalGroups.txt and leverage RDKit (Landrum, 2016) to detect the existence of the functional groups within a molecule. We leverage $3,300$ molecules from ChEBI-20 test split (Edwards et al., 2021), and adopt the query style as for large vision-language models (Li et al., 2023c), which queries the existence of the specific functional group in the molecule:
@@
Then, we detect whether the LGLM gives outputs meaning “Yes” or “No” following the practice in (Li et al., 2023c). For each molecule, we construct questions with positive answers for all kinds of functional groups detected in the molecule, and questions with negative answers for randomly sampled $6$ functional groups from the $38$ common functional groups in RDKit. Therefore, MotifHallu consists of $23,924$ query answer pairs. While it is easy to scale up MotifHallu by automatically considering more molecules and a broader scope of functional groups, we find that the current scale is already sufficient to demonstrate the hallucination phenomena in LGLMs.
@@ 4 Hierarchical Graph Tokenization
To mitigate the aforementioned issue, we propose a new strategy called HIerarchical GrapH Tokenization (HIGHT), which contains a hierarchical graph tokenizer that augments the input graph modality, as well as a hierarchical molecular instruction tuning dataset that augments the input language modality, to facilitate the alignment of molecule and language modalities.
@@ 4.1 Hierarchical Graph Tokenizer
Inspired by the success of hierarchical GNNs in self-supervised molecular representation learning (ZHANG et al., 2021; Zang et al., 2023), we transform the original molecular graph ${\mathcal{G}}$ into a hierarchical graph ${\mathcal{G}}^{\prime}$ with motif and graph nodes added in. Specifically, we leverage the Breaking of Retrosynthetically Interesting Chemical Substructures (BRICS) algorithm (Degen et al., 2008) to detect a set of $k$ motifs in ${\mathcal{G}}$, denoted as =${\mathcal{M}}=\{{\mathcal{M}}^{(1)},...,{\mathcal{M}}^{(k)},{\mathcal{M}}^{(k+1)}\}$, where ${\mathcal{M}}^{(k+1)}={\mathcal{G}}$ is the original molecule, without loss of generality. Furthermore, we denote the set of nodes and edges in ${\mathcal{M}}^{(i)}$ as ${\mathcal{V}}^{(i)}_{m}$ and ${\mathcal{E}}^{(i)}_{m}$, respectively. Then, we augment the original molecular graph ${\mathcal{G}}$ as ${\mathcal{G}}^{\prime}$ with augmented nodes ${\mathcal{V}}^{\prime}$ and edges ${\mathcal{E}}^{\prime}$:
@@
where $v^{(i)}_{m}$ is the motif super nodes added to the original molecule, and ${\mathcal{E}}_{ma}^{(i)}=\cup_{u\in{\mathcal{V}}_{m}^{(i)}}\{(u,v^{(i)}_{m})\}$ are the augmented edges connecting to the motif super node from nodes within the corresponding motif. We employ separate VQVAEs for atoms and motifs to learn meaningful code embeddings with several self-supervised learning tasks. The reconstructed attributes in Eq. 4 include atom types at the atom-level and the number of atoms at the motif-level, etc., following (Zang et al., 2023).
Meanwhile, merely feeding the motif tokens with node tokens to LLMs still can not help distinguish the motifs from nodes properly. Therefore, we propose to further attach positional encodings ${\bm{p}}$ to all of the tokens. We choose to use Laplacian positional embeddings (Dwivedi et al., 2020) while one could easily extend it with other variants (Ying et al., 2021). Meanwhile, as motif (and graph) tokens pose different semantic meanings from atom tokens, we adopt separate adapters for different types of tokens. Denote the motif tokens as ${\bm{h}}_{m}^{(i)}$ for motif ${\mathcal{M}}^{(i)}$, HIGHT tokenizer improves the LLM generation as:
$p_{\mathbf{\theta}}({\bm{a}}|{\bm{q}},{\bm{h}},{\bm{h}}_{m})=\prod_{i=1}^{l_{a}}p_{\mathbf{\theta}}({\bm{a}}_{i}|$ ${\bm{q}},f_{n}({\bm{h}}_{1}),...,f_{n}({\bm{h}}_{n}),$ (8)
$f_{m}({\bm{h}}_{m}^{(1)}),...,f_{m}({\bm{h}}_{m}^{(k)}),f_{g}({\bm{h}}_{m}^{(k+1)}),{\bm{a}}_{<i}),$
@@ 4.2 Hierarchical Graph Instruction Tuning Dataset
Although HIGHT tokenizer properly extracts the hierarchical information from the input graph modality, it remains challenging to properly align the language information to the corresponding graph information, without the appearance of the respective captions in the texts. For example, if the caption does not contain any information about the water solubility of the hydroxide functional group (“-OH”), LGLMs will never know that “-OH” motif corresponds to the water solubility of the molecule, despite that HIGHT tokenizer extracts the “-OH” token. In fact, the commonly used molecular instruction tuning curated from PubChem (Kim et al., 2022) in existing LGLMs (Liu et al., 2023d; Cao et al., 2023; Li et al., 2024), contains surprisingly little information about motifs. Some samples are given in Appendix B.2.
To this end, we propose HiPubChem, which augments the molecular instruction tuning dataset with captions of the functional groups. We consider both the positive and negative appearances of motifs when augmenting the instructions. For the positive case, we directly append the caption of all functional groups detected with RDKit:
This molecule has <#> of <functional group name> groups.
where <#> refers to the detected number of the functional group in the molecule, and <functional group name> refers to the name of the functional group as listed in Appendix B.2. For the negative case, we randomly sample $k_{\text{neg}}$ that do not appear in the molecule:
This molecule has no <functional group name> groups.
Despite the simple augmentation strategy, we find that HiPubChem significantly reduces the hallucination issue and improves the molecule-language alignment performance.
@@ 4.3 Hierarchical Graph Instruction Tuning
For the training of LGLM with HIGHT, we consider a simple two-stage instruction tuning as (Liu et al., 2023b; Cao et al., 2023).
Stage 1 Alignment Pretraining. We curate a new molecule-text paired dataset from PubChem following the pipeline of (Liu et al., 2023c). We set the cutoff date by Jan. 2024, and filter out unmatched pairs and low-quality data, which results in $295$k molecule-text pairs. Furthermore, we construct the HiPubChem-$295$k dataset based on the curated PubChem-$295$k dataset. The alignment pretraining stage mainly warms up the adapter to properly project the graph tokens with the LLM embedding space. To avoid feature distortion, both the LLM and the GNN encoder are frozen during stage 1.
Stage 2 Task-specific Instruction Tunning. With a properly trained adapter, we further leverage the task-specific instruction tuning datasets adapted from MoleculeNet (Wu et al., 2017), ChEBI-20 (Mendez et al., 2019), and Mol-Instructions (Fang et al., 2024). More details of the instruction tuning datasets are given in Appendix B. In Stage 2, we still keep the GNN encoder frozen, while tuning both the adapter and the LLM. The LLM is tuned using low-rank adaptation (i.e., LoRA) (Hu et al., 2022) following the common practice.
5 Experimental Evaluation
We conducted extensive experiments to evaluate HIGHT, comparing with previous node-centric tokenization, across $7$ real-world tasks including property prediction, molecular description, and chemical reaction prediction. We will briefly introduce the experimental setups, and leave more details in Appendix C.
5.1 Experimental Settings
Basically, we follow the common practice in the literature (Cao et al., 2023; Fang et al., 2024) to conduct our experiments.
Architecture. The GNN backbone used for producing graph tokens is a $5$-layer GIN (Xu et al., 2019) with a hidden dimension of $300$. The adapter is implemented as a single-layer MLP. The base LLM adopts the vicuna-v-1.3-7B (Chiang et al., 2023). The overall scale of parameters is around $6.9$B.
Baselines. We incorporate both the specialist molecular foundation/pretrained models, as well as LLM-based generalist models. The specialist models include expert models pretrained on large-scale molecular datasets and then finetuned on task-specific datasets such as KV-PLM (Zeng et al., 2022), GraphCL (You et al., 2020) and GraphMVP (Liu et al., 2022). The specialist models also include molecule-specialized foundation models that are trained with tremendous molecule-centric datasets such as MolT5-based methods (Edwards et al., 2022), Galactica (Taylor et al., 2022), MoMu (Su et al., 2022), MolFM (Luo et al., 2023a), Uni-Mol (Zhou et al., 2023), MolXPT (Liu et al., 2023f), GIT-Mol (Liu et al., 2024b), and BioMedGPT (Luo et al., 2023b). We adopt the results from previous works (Fang et al., 2024; Cao et al., 2023) when available.
For LLM-based generalist models, we consider LLMs such as ChatGPT (OpenAI, 2022), Llama (Touvron et al., 2023a) as well as instruction tuned LLMs such as Alpaca (Dubois et al., 2023), Baize (Xu et al., 2023), ChatGLM (Zeng et al., 2023) and Vicuna (Chiang et al., 2023). We also consider parameter-efficient finetuned LLMs using the backbone of Llama2 (Touvron et al., 2023b) as done by Mol-Instructions (Fang et al., 2024). For the node-centric based tokenization, we implement the baseline mainly based on InstructMol (Cao et al., 2023) with a VQVAE tokenizer from Mole-BERT (Xia et al., 2023). HIGHT is implemented based on the same architecture with only the tokenizer replaced. We use the suffix “-G” to refer to LLMs with only 2D graph input while using “-GS” to refer to LLMs with both 2D graph and 1D selfies input (Krenn et al., 2019; Fang et al., 2024; Cao et al., 2023). We do not include the baselines with “-GS” for some tasks as we find that incorporating the 1D input does not always bring improvements in the experiments.
Training and evaluation. We apply the same optimization protocol to tune LGLMs with node-centric and HIGHT tokenizers for fair comparisons. We train both models with stage 1 by $5$ epochs and stage 2 by $5$ to $50$ epochs as recommended by (Cao et al., 2023).
@@ 5.1 Experimental Settings
Table 1: Results of motif hallucinations on MotifHallu.

### v2

@@ 3.1 Node-Centric Tokenization
$p_{\mathbf{\theta}}({\bm{a}}|{\bm{q}},{\bm{h}})=\text{$\prod$}_{i=1}^{l_{a}}p_{\mathbf{\theta}}({\bm{a}}_{i}|{\bm{q}},f_{n}({\bm{h}}_{1}),...,f_{n}({\bm{h}}_{n}),{\bm{a}}_{<i}),$ (4)
where ${\bm{h}}_{1},...,{\bm{h}}_{n}$ are node embeddings from a GNN typically pretrained through self-supervised learning on large-scale molecular datasets such as ZINC250k (Sterling & Irwin, 2015), $f_{n}$ is the adapter to project the node tokens to the LLM tokens. There are various options to tokenize a molecule (Liu et al., 2023d). In this work, we consider a state-of-the-art tokenizer (Xia et al., 2023) that pretrains a VQVAE (van den Oord et al., 2017) with masked atoms modeling and constructs a codebook ${\mathcal{Z}}$ to discretize atoms: $z_{u}=\text{$\operatorname*{arg\,min}$}_{i}||{\bm{h}}_{u}-{\bm{e}}_{i}||_{2}$, where $z_{u}\in{\mathcal{Z}}$ is the quantized index of atom $u$, and ${\bm{e}}_{v}$ is the codebook embedding of the $i$-th entry. The codebook is trained through a reconstruction loss with respect to some attribute ${\bm{v}}_{i}$ of atom $i$:
${\mathcal{L}}_{r}=$ $\frac{1}{n}\sum_{i=1}^{n}(1-\frac{{\bm{v}}_{i}^{T}\hat{{\bm{v}}_{i}}}{||{\bm{v}}_{i}||\cdot||\hat{{\bm{v}}_{i}}||})^{\gamma}+\frac{1}{n}\sum_{i=1}^{n}||\text{sg}[{\bm{h}}_{i}]-{\bm{e}}_{z_{i}}||_{2}^{2}$ (5)
$+\frac{\beta}{2}\sum_{i=1}^{n}||\text{sg}[{\bm{e}}_{z_{i}}]-{\bm{h}}_{i}||_{2}^{2},$
where $\text{sg}[\cdot]$ is the stop-gradient operator in straight-through estimator (Bengio et al., 2013), $\hat{{\bm{v}}_{i}}$ is the reconstructed attribute of atom $i$ with a decoder, and $\beta$ is a hyperparamter. In Mole-BERT, the attribute is simply the type of atom. Mole-BERT also manually partitions the codebook into groups of common atoms such as carbon, nitrogen, and oxygen to avoid codebook conflicts (Xia et al., 2023).
@@
Figure 2: Illustration of hallucination caused by node-centric tokenization. With only node-level tokens, LLMs have to relate the nodes within a specific functional group to align useful molecular structures with the corresponding language descriptions. Yet, due to the arbitrary order of atoms and position biases in LLMs, it is hard to recognize each functional group, leading to severe hallucinations.
Intuitively, the trained atom tokens encode some contextual information, such as the neighbors of the atoms. However, node-centric tokenization makes the molecule-language alignment more challenging, as LLMs have to additionally relate the multiple nodes to align the corresponding texts during the instruction tuning process. Specifically, in molecules, motifs or functional groups usually capture rich semantics, and often share many common atoms such as carbon, nitrogen, and oxygen (Bohacek et al., 1996). As shown in Fig. 2, both the carboxylic acid (“R-COOH”) and the hydroperoxide (“R-OOH”) functional groups all contain two oxygen atoms and a hydrogen atom. For a molecule with hydroperoxide attached to a scaffold with carbon atoms, it would be hard for LLMs to distinguish which functional group is present in the molecule. Furthermore, due to the loss of positional information in the node-centric tokenization (Liang et al., 2023; Cao et al., 2023), the limited expressivity of GNNs (Xu et al., 2019) and the positional biases of auto-regressive LLMs (Lu et al., 2022), it is more challenging for LLMs to relate the desired nodes in a motif, which will lead to subpar molecule-language alignment.
@@ 3.2 Motif Hallucination
To understand the issue of node-centric tokenization more clearly, we construct a simple benchmark called MotifHallu, to concretize the hallucination of common functional groups by LGLMs. Specifically, we consider the $38$ common functional groups in RDKit [^1] https://github.com/rdkit/rdkit/blob/master/Data/FunctionalGroups.txt and leverage RDKit (Landrum, 2016) to detect the existence. We adopt $3,300$ molecules from ChEBI-20 (Edwards et al., 2021) and query the existence of a functional group:
@@
Then, we examine the outputs from LGLM meaning “Yes” or “No”. For each molecule, we construct questions with positive answers for all kinds of functional groups detected in the molecule, and questions with negative answers for randomly sampled $6$ functional groups from the remaining. Hence MotifHallu consists of $23,924$ questions. While it is easy to scale up MotifHallu with more molecules and functional groups, we find that the current scale is already sufficient to demonstrate the issue (Table 2).
@@ 4 Hierarchical Graph Tokenization
To improve the molecule-language alignment, we propose a new strategy called HIerarchical GrapH Tokenization (HIGHT), which contains a hierarchical graph tokenizer and a hierarchical molecular instruction tuning dataset to augment the inputs with hierarchical information.
@@ 4.1 Hierarchical Graph Tokenizer
Inspired by the success of hierarchical GNNs (Zhang et al., 2021; Zang et al., 2023), we transform the original molecular graph ${\mathcal{G}}$ into a hierarchical graph ${\mathcal{G}}^{\prime}$ with motif-level and molecule-level nodes added in. Specifically, we leverage the Breaking of Retrosynthetically Interesting Chemical Substructures (BRICS) algorithm (Degen et al., 2008) [^2] Note that HIGHT possesses a high degree of extensibility and can be augmented by incorporating advanced motif extraction techniques (such as (Zhang et al., 2021)). to detect and inject a set of $k+1$ supernodes, denoted as =${\mathcal{M}}=\{{\mathcal{M}}^{(1)},...,{\mathcal{M}}^{(k)},{\mathcal{M}}^{(k+1)}\}$, with $k$ motifs and the original molecule ${\mathcal{M}}^{(k+1)}={\mathcal{G}}$. Furthermore, denoting the set of nodes and edges in ${\mathcal{M}}^{(i)}$ as ${\mathcal{V}}^{(i)}_{m}$ and ${\mathcal{E}}^{(i)}_{m}$, respectively, we augment the original molecular graph ${\mathcal{G}}$ as ${\mathcal{G}}^{\prime}$ with augmented nodes ${\mathcal{V}}^{\prime}$ and edges ${\mathcal{E}}^{\prime}$:
@@
where $v^{(i)}_{m}$ is the motif super nodes added to the original molecule, and ${\mathcal{E}}_{ma}^{(i)}=\cup_{u\in{\mathcal{V}}_{m}^{(i)}}\{(u,v^{(i)}_{m})\}$ are the augmented edges connecting to the motif super node from nodes within the corresponding motif. We employ separate VQVAEs for atoms and motifs to learn meaningful code embeddings with several self-supervised learning tasks. The reconstructed attributes in Eq. 4 include atom types at the atom-level and the number of atoms at the motif-level (Zang et al., 2023).
Merely feeding the motif tokens with node tokens to LLMs still can not help distinguish the motifs from atoms properly, hence we propose to further attach positional encodings ${\bm{p}}$ to all of the tokens. We choose to use Laplacian positional embeddings (Dwivedi et al., 2020) while one could also adopt other variants (Ying et al., 2021). Since different types of tokens contain distinct semantics, we adopt separate adapters for different types of tokens. Denoting the motif tokens as ${\bm{h}}_{m}^{(i)}$ for motif ${\mathcal{M}}^{(i)}$, generation with HIGHT is:
$p_{\mathbf{\theta}}({\bm{a}}|{\bm{q}},{\bm{h}},{\bm{h}}_{m})=$ $_{i=1}^{l_{a}}p_{\mathbf{\theta}}({\bm{a}}_{i}|{\bm{q}},f_{n}({\bm{h}}_{1}),...,f_{n}({\bm{h}}_{n}),$ (7)
$f_{m}({\bm{h}}_{m}^{(1)}),...,f_{g}({\bm{h}}_{m}^{(k+1)}),{\bm{a}}_{<i}),$
@@ 4.2 Hierarchical Graph Instruction Tuning Dataset
Although HIGHT tokenizer properly extracts the hierarchical information from the input graph modality, it remains challenging to properly align the language to the corresponding molecular information, without the appearance of the respective captions in the texts. For example, if the caption does not contain any information about the water solubility of the hydroxide functional group (“-OH”), LGLMs will never know that “-OH” motif corresponds to the water solubility of the molecule, despite that HIGHT tokenizer extracts the “-OH” token. In fact, the commonly used molecular instruction tuning curated from PubChem (Kim et al., 2022) in existing LGLMs (Liu et al., 2023c; Cao et al., 2023; Li et al., 2024), contains surprisingly little information about motifs. Some samples are given in Appendix C.2.
To this end, we propose HiPubChem, which augments the molecular instruction tuning dataset with captions of the functional groups. We consider both the positive and negative appearances of motifs: For the positive case, we directly append the caption of all functional groups detected with RDKit. We also include a brief introduction of the functional groups to provide fine-grained information for molecule-language alignment. For the negative case, we randomly sample $k_{\text{neg}}$ motifs not appeared in the molecule to explicitly instruct LGLMs on the absence of the motifs. Despite the simple augmentation strategy, we find that HiPubChem significantly reduces the hallucination issue and improves the alignment performance.
@@ 4.3 Hierarchical Graph Instruction Tuning
We use a two-stage instruction tuning (Cao et al., 2023).
Stage 1 Alignment Pretraining. We curate a new molecule-text paired dataset from PubChem following the pipeline of Liu et al. (2023b). We set the cutoff date by Jan. 2024, and filter out unmatched pairs and low-quality data, which results in $295$k molecule-text pairs. Furthermore, we construct the HiPubChem-$295$k dataset. The first stage mainly warms up the adapter to properly project the graph tokens with the LLM embedding space. To avoid feature distortion, both the LLM and the GNN encoder are frozen.
Stage 2 Task-specific Instruction Tunning. With a properly trained adapter, we further leverage the task-specific instruction tuning datasets from MoleculeNet (Wu et al., 2017), ChEBI-20 (Mendez et al., 2019), and Mol-Instructions (Fang et al., 2024). More details are given in Appendix C. In Stage 2, we still keep the GNN encoder frozen, while tuning both the adapter and the LLM (with low-rank adaptation, i.e., LoRA (Hu et al., 2022)).
Table 1: Detailed results in motif hallucinations on MotifHallu. Due to the imbalance of samples from positive and negative classes, we incorporate diverse evaluation metrics to provide a detailed comparison between different methods in terms of hallucination.
Method Macro F1 $\uparrow$ F1 (pos) $\uparrow$ F1 (neg) $\uparrow$ Micro F1 $\uparrow$ AUROC $\uparrow$ Acc $\uparrow$ Yes Ratio
$4,124$ $19,800$
GIMLET (Zhao et al., 2023) 50.0 0.1 99.9 0.05 49.9 82.6 0.2
Galactica-6.7B (Taylor et al., 2022) 56.6 17.5 95.7 12.9 50.7 77.6 8.5
InstructMol (Cao et al., 2023) 52.6 95.7 9.5 28.3 48.4 20.0 94.5
HIGHT 66.8 85.5 48.2 29.7 53.2 39.1 69.4
Table 2: Results of motif hallucinations on MotifHallu.
@@
5 Experimental Evaluation
We conduct extensive experiments to compare HIGHT with previous node-centric tokenization across $14$ real-world tasks, including property prediction, molecular description, and chemical reaction prediction. The details and examples regarding the datasets and tasks involved in the experiments are given in Appendix C. We briefly introduce the setups below and leave the details in Appendix D.
5.1 Experimental settings
Architecture. The GNN backbone is a $5$-layer GIN (Xu et al., 2019) with a hidden dimension of $300$. The adapter is a single-layer MLP. We consider base LLMs of vicuna-v-1.3-7B (Chiang et al., 2023) for all the tasks and llama-2-7B-chat (Touvron et al., 2023b) for ablation studies.
Baselines. Since the focus of this work lies in the tokenization, our main comparison focuses on between HIGHT and node-centric tokenization. Nevertheless, we also include a series of existing LGLMs based on non-regression LLMs and regression LLMs, to provide an overview of the performance achieved by HIGHT. We would like to note that there are existing differences in pretraining data and information used between HIGHT and those baselines. For details, please refer to Table 7 in the Appendix.
For the node-centric based tokenization, we implement the baseline mainly based on InstructMol (Cao et al., 2023) with a VQVAE tokenizer from Mole-BERT (Xia et al., 2023). HIGHT is implemented based on the same architecture with only the tokenizer replaced. We use the suffix “-G” to refer to LGLMs with only 2D graph input and “-GS” to refer to LGLMs with both 2D graph and 1D selfies input (Krenn et al., 2019; Fang et al., 2024; Cao et al., 2023). We do not include the baselines with “-GS” for tasks other than MotifHallu as we find that incorporating the 1D input does not always bring improvements in the experiments.
For non-regression-based models, including the pretrained models such as KV-PLM (Zeng et al., 2022), GraphCL (You et al., 2020) and GraphMVP (Liu et al., 2022), and molecular foundation models that are trained with tremendous molecule-centric datasets such as MolT5-based methods (Edwards et al., 2022), Galactica (Taylor et al., 2022), MoMu (Su et al., 2022), MolFM (Luo et al., 2023a), Uni-Mol (Zhou et al., 2023), MolXPT (Liu et al., 2023e), GIT-Mol (Liu et al., 2024c), and BioMedGPT (Luo et al., 2023b). We adopt the results from the previous works Fang et al. (2024); Cao et al. (2023) if applicable.
For regression-based LGLMs, we consider LLMs such as ChatGPT (OpenAI, 2022), Llama (Touvron et al., 2023a) as well as instruction tuned LLMs such as Alpaca (Dubois et al., 2023), Baize (Xu et al., 2023), ChatGLM (Zeng et al., 2023) and Vicuna (Chiang et al., 2023). We also consider parameter-efficient finetuned LLMs using the backbone of llama2 (Touvron et al., 2023b) as done by Mol-Instructions (Fang et al., 2024).